OpenStudy (anonymous):

If f(x) = −2x − 4 and g(x) = x + 4, find (f over g)(−8). −3 3 −1 1

hero (hero):

f/g(-8) = f(-8)/g(-8) = (-2(8) - 4)/(-8 + 4) = (16 - 4)/(-4) = (12)/(-4) = -3

OpenStudy (anonymous):

thanx